# StyleDrive: Towards Driving-Style Aware Benchmarking of End-To-End Autonomous Driving

### v1

@@ Abstract
While personalization has been explored in traditional autonomous driving systems, it remains largely overlooked in end-to-end autonomous driving (E2EAD), despite its growing prominence. This gap is critical, as user-aligned behavior is essential for trust, comfort, and widespread adoption of autonomous vehicles. A core challenge is the lack of large-scale real-world datasets annotated with diverse and fine-grained driving preferences, hindering the development and evaluation of personalized E2EAD models. In this work, we present the first large-scale real-world dataset enriched with annotations capturing diverse driving preferences, establishing a foundation for personalization in E2EAD. We extract static environmental features from real-world road topology and infer dynamic contextual cues using a fine-tuned visual language model (VLM), enabling consistent and fine-grained scenario construction. Based on these scenarios, we derive objective preference annotations through behavioral distribution analysis and rule-based heuristics. To address the inherent subjectivity of driving style, we further employ the VLM to generate subjective annotations by jointly modeling scene semantics and driver behavior. Final high-quality labels are obtained through a human-in-the-loop verification process that fuses both perspectives. Building on this dataset, we propose the first benchmark for evaluating personalized E2EAD models. We assess several state-of-the-art models with and without preference conditioning, demonstrating that incorporating personalized preferences results in behavior more aligned with human driving. Our work lays the foundation for personalized E2EAD by providing a standardized platform to systematically integrate human preferences into data-driven E2EAD systems, catalyzing future research in human-centric autonomy.
@@ 1 Introduction
Personalization has long been recognized as a critical factor in enhancing user experience and building trust in autonomous driving systems Hasenjäger and Wersing (2017). In conventional modular pipelines—comprising components such as motion planning, trajectory prediction, and decision-making—personalized strategies have been widely studied, enabling systems to adapt to individual driving preferences. However, these methods often rely on isolated, scenario-specific adaptations Tian et al. (2022); Liu et al. (2024) or unreal human-in-the-loop (HuiL) simulation Ke et al. (2024), resulting in limited generalization across diverse and dynamic real-world contexts. Moreover, the fragmented nature of modular systems Zhu et al. (2018); Cui et al. (2024a); Kou et al. (2025) hinders scalability, as preference modeling is tightly coupled to specific modules and does not transfer easily to integrated learning systems. Due to these limitations, personalization remains largely underexplored in end-to-end autonomous driving (E2EAD), where perception, planning, and control are jointly optimized in a single architecture Chib and Singh (2023). This gap presents a significant barrier to realizing truly human-centric and adaptive autonomous driving at scale.
@@
This gap is particularly pressing. As E2EAD systems become increasingly capable and are deployed across diverse, open-world driving environments, aligning vehicle behavior with user preferences becomes essential for enhancing comfort, perceived safety, and long-term user acceptance Speidel et al. (2019); Aledhari et al. (2023). Yet integrating personalization into the end-to-end paradigm introduces unique challenges. Unlike modular systems operating within narrowly defined scenarios, E2EAD requires preference generalization across a broad spectrum of complex, real-world scenarios. Meeting this demand necessitates large-scale datasets that not only cover diverse traffic conditions but also provide fine-grained driving style annotations. Without such data, it is difficult to train or benchmark models that can scale with data and generalize across both users and environments. Consequently, data-driven personalization in E2EAD remains a critical and unresolved challenge.
To address this challenge, we introduce the first large-scale real-world dataset explicitly designed for personalized E2EAD. Our dataset captures both objective behavioral patterns and subjective driving preferences across a wide range of driving scenarios. Static environmental features are extracted from real-world road topologies, while dynamic context cues are inferred using a fine-tuned visual language model (VLM), enabling rich semantic scene understanding. Based on these scenarios, we generate objective annotations through behavior distribution analysis and rule-based heuristics. To incorporate subjectivity, we further employ the VLM to synthesize preference labels that integrate scene semantics and inferred driver intent, followed by human-in-the-loop verification to ensure quality and consistency. Building upon this dataset, we establish the first benchmark for evaluating personalized E2EAD models. We conduct extensive experiments on multiple state-of-the-art architectures, with and without preference conditioning, and show that incorporating personalization significantly improves alignment with human-like driving behavior.
Our work aims to bridge the gap between the growing momentum in E2EAD and the longstanding need for human-centric design in autonomous driving systems. By introducing both a large-scale data foundation and a standardized evaluation benchmark, we seek to facilitate future research on personalized, adaptive, and trustworthy autonomous driving. In summary, our key contributions are:
@@
  • A multi-stage annotation pipeline that integrates behavior analysis, rule-based heuristics, visual language models (VLMs), and human-in-the-loop verification to generate consistent and high-quality preference labels.
  • The first benchmark for personalized E2EAD, which enables standardized, quantitative comparison of preference-conditioned behavior across different model architectures.
@@
Abbreviations: OL = Open-Loop, CL = Closed-Loop, Semi-CL = Semi-Closed-Loop, HuiL = Human-in-the-Loop, E2E = End-to-End driving model support, Style = With Style or Preference Annotation
@@
HuiL-RM Li et al. (2023) HuiL Ramp-Merge OL
HuiL-CF Zhao et al. (2022) HuiL Car-Following OL
HuiL-LC Liao et al. (2023) HuiL Lane-Change OL
HuiL-Mul Ke et al. (2024) HuiL City OL
@@ 2 Related Work
Personalization has long been recognized as a key factor in enhancing user comfort, trust, and long-term acceptance of autonomous vehicles Liao et al. (2024). In conventional modular pipelines—comprising separate stages for perception, prediction, and planning—personalized strategies have been extensively explored. These methods allow the system to adapt to user-specific preferences in specific driving scenarios, such as desired following distance, acceleration aggressiveness, and lane-change behavior Müller et al. (2018); Huang et al. (2021); Tian et al. (2022); Rosero et al. (2024); Liu et al. (2024).
Recently, the emergence of large language models (LLMs) has opened new possibilities for personalization in autonomous driving. Several studies have explored leveraging LLMs to encode user intent, interpret high-level preferences, and influence downstream decision-making Cui et al. (2024a); Xu et al. (2024); Cui et al. (2024b); Kou et al. (2025); Lin et al. (2024). However, these approaches typically operate within modular or semi-modular frameworks and often rely on handcrafted integration points or high-level perception outputs. MAVERIC Schrum et al. (2024) represents a notable early attempt to enable end-to-end personalized driving via a unified latent representation of user style. Nevertheless, it is developed in a semi-simulated environment with limited sensory inputs—excluding modalities like LiDAR—and lacks a standardized benchmark for comparing methods, which constrains its scalability and broader applicability in real-world settings.
Despite these advancements, personalized E2EAD remains largely underexplored. Existing research seldom addresses how user preferences can be directly embedded into unified end-to-end frameworks. A key bottleneck in this direction is the lack of large-scale, real-world datasets with fine-grained annotations of personalized driving styles.
@@
End-to-end autonomous driving (E2EAD) has emerged as a promising paradigm that maps raw sensor inputs directly to driving actions or planned trajectories using deep neural networks Chen et al. (2024a). Early work Codevilla et al. (2018); Kendall et al. (2019); Chen et al. (2020a) demonstrated the feasibility of E2E systems through imitation and reinforcement learning in both real-world and simulated settings. Over the past decade, research has advanced toward more sophisticated architectures incorporating temporal reasoning, multimodal fusion, and Transformer-based backbones Prakash et al. (2021); Jia et al. (2023); Shao et al. (2023). Recent efforts have also focused on data-efficient training and task-conditioned control across diverse scenarios Kong et al. (2025).
Despite this progress, most E2EAD models are optimized for average-case behavior, lacking mechanisms to account for driver-specific preferences or personalized objectives. While recent architectures improve generalization and interpretability Chen et al. (2020b); Li et al. (2024); Zheng et al. (2024); Chen et al. (2024b); Jia et al. (2025), they remain limited in capturing nuanced, individualized driving styles. Moreover, the absence of dedicated datasets and standardized benchmarks for evaluating personalization continues to hinder systematic advancement in this area.
To address these limitations, we build on recent advances in model capacity and controllability to develop and evaluate E2EAD systems that explicitly incorporate user driving preferences. Our work represents the first systematic attempt to align end-to-end driving behavior with real-world personalized objectives, paving the way for more adaptive and user-aligned autonomous vehicles.
@@
Benchmarking is essential for evaluating autonomous driving models. Early E2EAD benchmarks such as nuScenes Caesar et al. (2020) and OpenScene Peng et al. (2023) adopt open-loop protocols using real-world data, but lack real-time interaction and fail to capture feedback effects or behavioral consequences. Recent works have introduced closed-loop benchmarks in simulation, including Longest6 Chitta et al. (2022), MetaDrive Li et al. (2022), CARLA Leaderboard V2 Contributors (2024), and Bench2Drive Jia et al. (2024). These allow online policy evaluation but rely on synthetic sensor data and simplified environments. Semi-simulated platforms like NAVSIM Dauner et al. (2024) strike a balance by replaying real-world scenes in simulation. However, existing benchmarks remain focused on task performance and ignore user preferences, offering no means to evaluate personalized driving behavior.
Efforts to benchmark personalized autonomous driving have resulted in two main types of datasets: human-in-the-loop (HuiL) simulation environments and real-world driving records. HuiL-based datasets are typically collected through controlled interactions between human drivers and simulated agents in targeted scenarios such as ramp merging (HuiL-RM) Li et al. (2023), car following (HuiL-CF) Zhao et al. (2022), lane changing (HuiL-LC) Liao et al. (2023), and multiple scenarios (HuiL-Mul) Ke et al. (2024). While these setups enable precise behavior capture and fine-grained preference modeling, they often rely on simplified simulation engines and handcrafted environments, resulting in a significant domain gap from real-world conditions. In contrast, real-world datasets such as UAH-DriveSet Romera et al. (2016), Brain4Cars Jain et al. (2016), and PDB Wei et al. (2025) provide behavioral annotations related to aggressiveness, drowsiness, or maneuver intent. These datasets have contributed substantially to understanding driver behavior and intent modeling. However, they are typically constrained to narrow driving contexts—such as highways or intersections—and lack the visual, semantic, and behavioral diversity needed for general-purpose deployment. More importantly, they are not structured for end-to-end learning: most only provide high-level behavior labels without continuous perception-to-control supervision, and none include standardized evaluation protocols for comparing personalized driving policies.
To bridge the gap between these two lines of work, we introduce the first benchmark tailored for personalized E2EAD. Built on real-world data, our benchmark supports both supervised learning and semi-closed-loop evaluation of preference-conditioned driving behavior. It provides structured annotations of user driving styles, integrates them into a unified end-to-end learning pipeline, and enables systematic, reproducible evaluation of personalized E2E policies across diverse scenarios.
@@ 3 StyleDrive Dataset
The StyleDrive dataset is constructed upon the large-scale real-world autonomous driving dataset OpenScene, and contains nearly 30,000 driving scenarios annotated with style preferences. In addition to the original OpenScene labels, we enrich the dataset with broader road topology annotations and propose a unified framework for modeling and labeling personalized driving preferences.
@@ 3.1 Framework for Modeling and Annotation of Driving Preference
To enable reliable and interpretable driving style analysis, we construct a hierarchical modeling and annotation framework, as shown in Figure 2. We extract static environmental features from real-world road topology and infer dynamic contextual cues using a fine-tuned visual language model (VLM), enabling consistent and fine-grained scenario construction. Based on these scenarios, we derive objective preference annotations through behavioral distribution analysis and rule-based heuristics. To address the inherent subjectivity of driving style, we further employ the VLM to generate subjective annotations by jointly modeling scene semantics and driver behavior. Final high-quality labels are obtained through a human-in-the-loop verification process that fuses both perspectives.
@@
To facilitate structured analysis of driving behavior, we first categorize each driving clip into a coarse-grained traffic scenario type based on the topology of high-definition (HD) maps. This classification establishes a physically grounded and interpretable basis for subsequent semantic modeling. Specifically, we begin by enriching the topological structures based on the original nuPlan HD maps (details in Sect. VIII). We then align each clip data with detailed map elements—including lane geometries, stop lines, crosswalks, and intersection layouts—to assign each segment to one of 11 primary traffic scenario types. Each type captures the static constraints and navigational features of the road environment, providing a foundational layer of context for interpreting driving behavior. The complete type list is shown in the first column of Tab. 2.
@@
To enrich the traffic scenario classification with dynamic, context-sensitive information, we apply a fine-tuned vision-language model (VLM) to extract high-level semantic cues from driving videos. While generic VLMs lack task-specific perception for autonomous driving, we adapt the Video-LLaMA3 model Zhang et al. (2025) via lightweight fine-tuning using the LingoQA dataset - a multimodal dataset for reasoning about road semantics (more details in Sect. X).
@@
For each traffic scenario type, we craft targeted prompts related to key driving events. These include questions such as whether a lead car is present in the same lane, whether the ego is merging, or whether pedestrians are visible in crosswalk zones. The model’s responses are parsed into structured semantic attributes, such as presence of a lead vehicle, merge risk, pedestrian involvement, and turning direction. This dynamic semantic layer augments the static topology with interaction-aware cues, supporting downstream preference inference through both rule-based and VLM-driven reasoning.
Objective Preference Annotation via Rule-Based Heuristics

### v2

@@ Abstract
Personalization, while extensively studied in conventional autonomous driving pipelines, has been largely overlooked in the context of end-to-end autonomous driving (E2EAD), despite its critical role in fostering user trust, safety perception, and real-world adoption. A primary bottleneck is the absence of large-scale real-world datasets that systematically capture driving preferences, severely limiting the development and evaluation of personalized E2EAD models. In this work, we introduce the first large-scale real-world dataset explicitly curated for personalized E2EAD, integrating comprehensive scene topology with rich dynamic context derived from agent dynamics and semantics inferred via a fine-tuned vision-language model (VLM). We propose a hybrid annotation pipeline that combines behavioral analysis, rule-and-distribution-based heuristics, and subjective semantic modeling guided by VLM reasoning, with final refinement through human-in-the-loop verification. Building upon this dataset, we introduce the first standardized benchmark for systematically evaluating personalized E2EAD models. Empirical evaluations on state-of-the-art architectures demonstrate that incorporating personalized driving preferences significantly improves behavioral alignment with human demonstrations.
@@ 1 Introduction
As autonomous driving (AD) technologies continue to mature, one of the key factors for translating these systems into real-world products is personalization. Tailoring vehicle behavior to individual user preferences is essential for enhancing user experience, building trust, and fostering long-term adoption Hasenjäger and Wersing (2017). Traditional modular systems have extensively explored personalized strategies, enabling adaptations to specific driving preferences. However, these methods often rely on isolated, scenario-specific adaptations Tian et al. (2022) or unrealistic human-in-the-loop (HITL) simulation Ke et al. (2024), which severely limit their generalization to dynamic, real-world environments. Furthermore, the fragmented nature of modular systems Zhu et al. (2018); Cui et al. (2024); Kou et al. (2025) hinders scalability to massive real-world data. Due to these limitations, personalization remains largely underexplored in end-to-end autonomous driving (E2EAD), where perception, planning, and control are integrated in an unified architecture Codevilla et al. (2018). This gap presents a significant barrier to realizing human-centric AD at scale.
@@
This gap is particularly pressing. As E2EAD systems become increasingly capable and are deployed across diverse, open-world driving environments, aligning vehicle behavior with user preferences becomes essential for enhancing comfort, perceived safety, and long-term user acceptance Speidel et al. (2019); Aledhari et al. (2023). Yet integrating personalization into the end-to-end paradigm introduces unique challenges. Unlike modular systems operating within narrowly defined scenarios, E2EAD requires preference generalization across a broad spectrum of complex, real-world scenarios. Meeting this demand necessitates large-scale datasets that not only cover diverse traffic conditions but also provide well-crafted driving style annotations.
To fill this critical gap, we introduce the first large-scale real-world dataset purposefully designed for personalized E2EAD research. Our dataset captures both objective behavioral patterns and subjective driving preferences across a wide range of driving scenarios. Static environmental features are firstly extracted from real-world road topologies, while dynamic context cues are inferred using a fine-tuned visual language model (VLM), enabling rich semantic scene understanding. Based on these features, we generate objective annotations through behavior distribution analysis and rule&distribution-based heuristics. To incorporate subjectivity, we further employ the VLM to label preferences considering static and dynamic scene semantics. The pipeline ends with a label fusion mechanism to ensure quality and consistency. Building upon this dataset, we establish the first benchmark for evaluating personalized E2EAD models. We conduct extensive experiments on multiple state-of-the-art architectures, with and without preference conditioning, and show that incorporating personalization significantly improves alignment with human-like driving behavior.
The motivation and overview of our work are illustrated in Fig. 1, aiming to bridge the gap between the growing momentum in E2EAD and the longstanding need for personalization in AD. In summary, our key contributions are:
@@
  • A multi-stage annotation pipeline that integrates behavior feature analysis, rule&distribution-based heuristics, VLM reasoning, and human-in-the-loop validation to generate high-quality well-crafted preference labels.
  • The first benchmark for personalized E2EAD, which enables standardized, quantitative comparison of preference-conditioned behavior across four different SOTA model architectures.
@@
Abbreviations: OL = Open-Loop, CL = Closed-Loop, Semi-CL = Semi-Closed-Loop, HITL = Human-in-the-Loop, E2E = End-to-End driving model support, Style = With Style or Preference Annotation
@@
HITL-RampMerging Li et al. (2023) HITL Ramp-Merge OL
HITL-CarFollowing Zhao et al. (2022) HITL Car-Following OL
HITL-LaneChange Liao et al. (2023) HITL Lane-Change OL
HITL-MultiScene Ke et al. (2024) HITL City OL
@@ 2 Related Work
Personalization has long been recognized as a key factor to enhance user comfort, trust, and acceptance of AD Liao et al. (2024). In conventional modular pipelines, personalized strategies have been extensively explored, where the system can adapt to users’ preferences in stand-alone driving scenarios, such as car following, ramp merge, and lane-change behaviors Zhao et al. (2022); Li et al. (2023); Liao et al. (2023). Other early attempts personalize models in unrealistic human-in-the-loop (HITL) simulation Ke et al. (2024). However, these approaches often suffer from poor generalization in dynamic, real-world environments.
Recent advances in large language models (LLMs) have introduced new possibilities for personalization in AD, enabling the encoding of user intent and preferences to influence downstream decision-making Cui et al. (2024); Xu et al. (2024); Kou et al. (2025). Despite their potential, these methods remain constrained by modular frameworks and often depend on handcrafted features. MAVERIC Schrum et al. (2024) represents a notable early attempt to enable end-to-end personalized driving. Nevertheless, in the absence of standardized datasets and benchmarks, these methods face challenges in reproducibility and scalability. Despite these advancements, existing work seldom addresses how to merge user preferences in an E2EAD manner. A critical bottleneck is the lack of large-scale, real-world datasets with well-crafted annotations of personalized driving styles.
@@
End-to-end autonomous driving represents a promising paradigm that maps raw sensor inputs directly to driving actions or planned trajectories. Early work Codevilla et al. (2018) demonstrated the feasibility of E2EAD. Over the past decade, research has advanced toward more sophisticated architectures incorporating temporal reasoning, multimodal fusion, and Transformer-based methods Prakash et al. (2021); Jia et al. (2023); Shao et al. (2023).
Despite this progress, most E2EAD models are optimized for average-level objectives, lacking the capacity to adapt to user-specific preferences. While recent architectures improve generalization and interpretability Li et al. (2024); Zheng et al. (2024); Jia et al. (2025), they remain limited in capturing individualized driving styles. Moreover, the absence of dedicated datasets and standardized benchmarks for evaluating personalization continues to hinder systematic advancement in this area. To address these limitations, we propose a framework for learning style-conditioned E2EAD models and establish a benchmark to evaluate their alignment with human driving behavior.
@@
Benchmarking is fundamental to the development and evaluation of AD models. Typical AD datasets and benchmarks are summarized in Tab. 1. Early E2EAD benchmarks Caesar et al. (2020); Peng et al. (2023) adopt non-interactive open-loop settings with real-world data, but fail to capture behavioral feedback. Recent works have introduced closed-loop benchmarks in simulation, such as Longest6 Chitta et al. (2022), MetaDrive Li et al. (2022), CARLA Series Contributors (2024), which allow online policy evaluation but rely entirely on simulated scenarios. Semi-simulated platforms like NAVSIM Dauner et al. (2024) strike a balance by simulating behavior feedback based on real-world scenes. However, existing benchmarks remain focused on task performance and ignore user preferences, cannot evaluate personalized driving behavior.
Efforts to benchmark personalized AD have led to two main dataset categories: human-in-the-loop (HITL) simulations and real-world driving logs. 1) HITL datasets are typically collected in controlled scenarios such as ramp merging Li et al. (2023), car following Zhao et al. (2022), lane changing Liao et al. (2023), and multiple scenarios Ke et al. (2024). While these setups enable preference modeling, they often rely on simplified simulation engines and curated environments, introducing a domain gap between real-world conditions and unrealistic scenarios. 2) Real-world datasets such as UAH-DriveSet Romera et al. (2016), Brain4Cars Jain et al. (2016), and PDB Wei et al. (2025) provide behavioral annotations related to diving intent. These datasets have contributed substantially to understanding driver behavior and intent modeling. However, they are typically constrained to narrow driving contexts, such as highways or intersections. More importantly, they are not structured for end-to-end learning: most only provide high-level behavior labels without continuous perception-to-control supervision, and none include standardized evaluation protocols for comparing personalized driving policies.
To bridge the gap between E2EAD and personalized AD, we introduce the first benchmark tailored for personalized E2EAD. Built on real-world data, our benchmark supports both supervised learning and semi-closed-loop evaluation of preference-conditioned policies, enabling reproducible evaluation of personalized E2E across diverse scenarios.
@@ 3 StyleDrive Dataset
The StyleDrive dataset is sampled and constructed from the large-scale AD dataset OpenScene Peng et al. (2023), and comprises nearly 30k driving scenarios annotated with style preferences. OpenScene features trajectories collected from over 16 drivers across diverse urban and suburban environments in both Singapore and the USA. This diversity offers a rich foundation for labeling user-specific tendencies in various scenarios (also illustrated in Fig. 4). Beyond the original OpenScene data, we introduce a unified framework for modeling and annotating personalized driving preferences.
@@ 3.1 Framework for Modeling and Annotation of Driving Preference
Framework Overview.
To enable reliable and interpretable driving style analysis, we construct a hierarchical modeling and annotation framework, as shown in Fig. 2. We extract static environmental features from real-world road topology and dynamic motion features from driving logs, and infer dynamic contextual cues using a fine-tuned visual language model (VLM), enabling consistent and fine-grained scenario construction. From these constructed scenario features, we derive objective preference annotations via behavioral distribution analysis and rule-based heuristics. To capture the inherent subjectivity of driving style, we further leverage the VLM to generate subjective annotations by jointly modeling scene semantics and driver behavior. Final high-quality labels are obtained through a human-in-the-loop verification process that consolidates both perspectives.
@@
To facilitate structured analysis of driving behavior, we first categorize each driving clip into a coarse-grained traffic scenario type based on the topology of high-definition (HD) maps. This categorization provides a physically grounded and semantically interpretable foundation for subsequent scenario modeling. Specifically, we begin by enriching the topological structures based on the original nuPlan HD maps (details in Sect.3 of the Supplementary). We then align each clip data with detailed map elements, such as lane geometries, stop lines, crosswalks, and intersection layouts, to assign it to one of eleven primary traffic scenario types. Each type captures the static constraints and navigational features of the road environment, serving as a structural prior to interpreting the downstream driving behavior. The complete type list is shown in the left part of Fig. 3.
@@
To enrich the traffic scenario classification with dynamic, context-sensitive information, we apply a fine-tuned vision-language model (VLM) to extract high-level semantic cues from driving videos. Since generic VLMs lack task-specific perception and reasoning capabilities for AD, we adapt the Video-LLaMA3 model Zhang et al. (2025) via lightweight fine-tuning using the LingoQA dataset Marcu et al. (2024) - a multimodal dataset for reasoning about road semantics (more details in Sect.5 of the Supplementary).
@@
For each traffic scenario type, we craft targeted prompts related to key driving events. These include questions such as whether a lead car is present in the same lane, whether the ego is merging, or whether pedestrians are visible in crosswalk zones. The model responses are then parsed into structured semantic attributes, such as the presence of lead vehicles, merging behavior, pedestrian involvement, and turning intent. This dynamic semantic layer augments the static topology with interaction-aware, temporally grounded cues, thereby enhancing downstream preference modeling via both rule&distribution-based and VLM-based reasoning.
Objective Preference Annotation via Rule&Distribution-based Heuristics